OpenStudy (kittyloli):

Line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to Line AB? 5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0

sam (.sam.):

Any ideas?

OpenStudy (kittyloli):

I think it might be B? But I'm really bad at geometry so I don't know... I need to finish this class for graduation... And this is my second year taking Geometry... Math isn't my strong suit

sam (.sam.):

You're in the second year of class already, you should have a grasp of this question, and the question is very easy

OpenStudy (kittyloli):

I barely passed the first year...

sam (.sam.):

You just start by finding the gradient first, and then, since the equation of the line that passes through the origin will have a coordinate of (0,0), you will then use \((y-y_1)=m(x-x_1)\) to construct your new equation.

OpenStudy (kittyloli):

Thank you

OpenStudy (dustin_whitelock):

@Kittyloli Use this point-slope form I told you about earlier, instead use this two-point slope form \[(y-y1)\div(x-x1) = (y1-y2)\div(x1-x2)\] Put the coordinates of A and B in x1, y1, x2, y2 resp. Now you got the equation of your line Arrange this in the form of ax+by=k (Note: k may be zero) Now, k can be any real number, and you get your answer :)

OpenStudy (kittyloli):

I got it :3